OpenStudy (anonymous):

A population of squirrels numbered 50 on 30 June 1980. On 30 June 1990, the population of squirrels numbered 600 000 000. The population can be modelled using the exponential population model.

OpenStudy (anonymous):

could you do the last one elizated?

OpenStudy (anonymous):

Determine the yearly growth rate of this population of squirrels. Estimate the squirrel population on 30 June 1985

OpenStudy (anonymous):

Sorry Paul, I'm not good with exponential functions. I like my lines straight.

OpenStudy (anonymous):

i should just put a picture of some girl on my profile and change my name to get some help around here... gah

OpenStudy (anonymous):

N = P * r^t N = population, r = rate of growth t = time P = initial population

OpenStudy (anonymous):

so all i need to do is know how to use logs...

OpenStudy (anonymous):

yup, that would be helpful

OpenStudy (anonymous):

eliza do you know?

OpenStudy (anonymous):

im having a bit of trouble getting r from \[10 \log (r) = \log 12*10^7\]

OpenStudy (anonymous):

wait up

OpenStudy (anonymous):

no. i don't know what log is, don't expect me to know how to use it!

OpenStudy (anonymous):

i think i have it

OpenStudy (anonymous):

log r = .70791...

OpenStudy (anonymous):

ahhh im so close aren't i

OpenStudy (anonymous):

brain ... growing

OpenStudy (anonymous):

yeah. try using google

OpenStudy (anonymous):

must figure out by self

OpenStudy (anonymous):

or wolfram alpha, both have built in calculators

OpenStudy (anonymous):

i feel like theres some simple step im missing that i should know... can you tell me it?